# Muon g-2 and lepton flavor violation in supersymmetric GUTs

### v1

@@ 4.1 Slepton masses and LFV
Figure 4: BR($\mu\rightarrow e\gamma$) vs $m_{\chi}$ for the three models discussed in the text. Order of the panels, symbols and color codes are as in Fig. 3. In addition the points with relic density inside the Plank bounds are colored in green.
Figure 5: GUT scale values of $m_{L}$ vesus $m_{R}$. Order of panels, symbols and colors are as in Fig. 3. We mark in green the points with relic density inside Planck bounds all the points satisfy the BR($\mu\rightarrow e\gamma$) constraint.
@@
Another important observation is the correlation of the BR($\mu\rightarrow e\gamma$) with other cLFV decays, such as $\tau\rightarrow\mu\gamma$. Generic see-saw models predict more or less the same rate for these decays, while the experimental bounds differ by several orders of magnitude. On the other hand, in general, the prediction for flavor mixing related to the family symmetry can allow a different hierarchy for the LFV $\tau$ and $\mu$ decays. In Fig. 6 we display BR($\mu\rightarrow e\gamma$) vs BR($\tau\rightarrow\mu\gamma$), and observe that scenarios with family symmetries can predict branching ratios for these decays in ranges closer to the experimental values. In addition to fitting the models with the current cLFV constraints, the solutions will be potentially tested soon by the improvements and upgrades in MEG II MEGII:2018kmf ; MEGII:2023ltw ; MEGII:2023fog , Mu2e Hedges:2022tnh , COMET COMET:2018auw and DeeMe Teshima:2019orf . decays
@@ 4.2 Muon g-2 and SUSY Masses
We have also displayed the scales for $M_{3}$ in Fig. 8 which summarizes the statements of suppressed flavor mixings in the quark sector due to the heavy gluinos in the regions favored by the muon $g-2$ condition. Recall that despite the light scales for $m_{L}$ and $m_{R}$, the dominant enhancement in the squark masses arises mostly from $M_{3}$ through RGEs. The large squark masses due to the heavy gluinos are also helpful in realizing the desired SM-like Higgs boson mass. In the MSSM framework the loop correction to the SM-like Higgs boson mass is given by Carena:2012mw
@@
Figure 11: Spin-independent cross section in correlation with the LSP mass. Order of panels, symbols and colors are as in Figs. 3 and 5. The solid lines are bounds from XENONnT XENON:2023cxc (black) and LZ LZ:2022lsv (red), while dashed lines are the projected limits from XENON-nT XENON:2015gkh (black) and DARWIN DARWIN:2016hyl (blue).
@@ 4.3 LFV versus LHC spectroscopy
Figure 12: $m_{\tilde{\tau}}-m_{\chi}$ plots for models with BR($\mu\rightarrow e\gamma$) below the current bounds. Order of panels, symbols and colors are as in Figs. 3 and 5. The solid lines indicate the contours extracted by the LHC searches, while the cyan marks highlight models with BR($\mu\rightarrow e\gamma$) one order of magnitude below the current bound.
@@
Figure 13: $m_{\tilde{\chi}^{\pm}}-m_{\chi}$ plots for models with BR($\mu\rightarrow e\gamma$) below the current bounds. Order of panels, symbols and colors are as in Figs. 3 and 5. The solid curves indicate the contours extracted by the LHC searches, while the cyan marks highlight models with BR($\mu\rightarrow e\gamma$) one order of magnitude below the current bound.
@@
$BR(\tau\rightarrow\mu\gamma)$ $4.8\cdot 10^{-14}$ $3.1\cdot 10^{-13}$ $1.1\cdot 10^{-12}$ $1.2\cdot 10^{-142}$
@@
Finally, we display two tables of benchmark points that are representative of our findings. The points listed in Table 2 belong to the seesaw models, and those in Table 3 represent the solutions realized in models with $\epsilon=0.05$ and $\epsilon=0.2$, as indicated in the header of Table 3. All of the benchmark points of seesaw scenario presented in Table 2 are compatible with the muon $g-2$ solutions within $2\sigma$, the Planck bound on the LSP neutralino relic density, and the constraint from $\mu\rightarrow e\gamma$. In addition, all these points, including P1, remain outside the contours of the current LHC bounds on the chargino and stau masses.

### v3

@@ 4.1 Slepton masses and LFV
Refer to caption Refer to caption Refer to caption
a
Figure 4: BR($\mu\rightarrow e\gamma$) vs $m_{\chi}$ for the three models discussed in the text. Order of the panels, symbols and color codes are as in Fig. 3. In all the models the prediction for $\Omega h^{2}$ is below the upper limit of eq. Planck:2015fie , the subset of points that are also above the lower bound are marked in green.
Refer to caption Refer to caption Refer to caption
Figure 5: GUT scale values of $m_{L}$ vesus $m_{R}$. Order of panels, symbols and colors are as in Fig. 3. All the points satisfy the BR($\mu\rightarrow e\gamma$) constraint. In all the models the prediction for $\Omega h^{2}$ is below the upper limit of eq. Planck:2015fie , the subset of points that are also above the lower bound are marked in green.
@@
Refer to caption Refer to caption Refer to caption
@@
Another important observation is the correlation of the BR($\mu\rightarrow e\gamma$) with other cLFV decays, such as $\tau\rightarrow\mu\gamma$. Generic see-saw models predict more or less the same rate for these decays, while the experimental bounds differ by several orders of magnitude. On the other hand, in general, the prediction for flavor mixing related to the family symmetry can allow a different hierarchy for the LFV $\tau$ and $\mu$ decays. In Fig. 6 we display BR($\mu\rightarrow e\gamma$) vs BR($\tau\rightarrow\mu\gamma$), and observe that scenarios with family symmetries can predict branching ratios for these decays in ranges closer to the experimental values. In addition to fitting the models with the current cLFV constraints, the solutions will be potentially tested soon by the improvements and upgrades in MEG II MEGII:2018kmf ; MEGII:2023ltw ; MEGII:2023fog , Mu2e Hedges:2022tnh , COMET COMET:2018auw and DeeMe Teshima:2019orf .
@@ 4.2 Muon $g-2$ and SUSY Masses
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
We have also displayed the scales for $M_{3}$ in Fig. 8 which summarizes the statements of suppressed flavor mixings in the quark sector due to the heavy gluinos in the regions favored by the muon $g-2$ condition. Recall that despite the light scales for $m_{L}$ and $m_{R}$, the dominant enhancement in the squark masses arises mostly from $M_{3}$ through RGEs. The large squark masses due to the heavy gluinos are also helpful in realizing the desired SM-like Higgs boson mass. To compute the Higgs boson mass we follow the procedure in Staub:2017jnp , as implemented in SPheno. However, in our discussion it is useful to use the approximate loop correction to the SM-like Higgs boson mass given in Carena:2012mw ,
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 11: Spin-independent cross section in correlation with the LSP mass. Order of panels, symbols and colors are as in Figs. 3 and 5. The solid lines are bounds from XENONnT XENON:2023cxc (black) and LZ LZ:2022lsv (red), while dashed lines are the projected limits from XENON-nT XENON:2015gkh (black) and DARWIN DARWIN:2016hyl (blue). The points of the upper panels predict neutralino relic density below the upper bound in eq. 25, while the subset of points contained in the lower panels are also above the lower bound in eq. 25.
@@
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 12: $m_{\tilde{\tau}}-m_{\chi}$ plots for models with BR($\mu\rightarrow e\gamma$) below the current bounds. Order of panels, symbols and colors are as in Figs. 3 and 5. The solid lines indicate the contours extracted by the LHC searches, while the cyan marks highlight models with BR($\mu\rightarrow e\gamma$) one order of magnitude below the current bound. The points of the upper panels predict neutralino relic density below the upper bound in eq. 25, while the subset of points contained in the lower panels are also above the lower bound in eq. 25.
@@ 4.3 LFV versus LHC spectroscopy
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 13: $m_{\tilde{\chi}^{\pm}}-m_{\chi}$ plots for models with BR($\mu\rightarrow e\gamma$) below the current bounds. Order of panels, symbols and colors are as in Figs. 3 and 5. The solid curves indicate the contours extracted by the LHC searches, while the cyan marks highlight models with BR($\mu\rightarrow e\gamma$) one order of magnitude below the current bound. The points of the upper panels predict neutralino relic density below the upper bound in eq. 25, while the subset of points contained in the lower panels are also above the lower bound in eq. 25.
@@
$BR(\tau\rightarrow\mu\gamma)$ $4.8\cdot 10^{-14}$ $3.1\cdot 10^{-13}$ $1.1\cdot 10^{-12}$ $1.2\cdot 10^{-14}$
@@
In order to make an explicit summary of the impact of LFV on signals from the MSSM models derived from $4-2-2$, we chose some representative points in the allowed solutions displayed in Figs. 12 and 13. These points can provide an indication of the magintude of the SUSY particles in models that satisfy $g-2$ due to large gluino contributions, the implication of DM in the prediction of compress SUSY spectrum and the relevance of alternative experimental signals to complement the LHC SUSY searches. Finally, we display two tables of benchmark points that are representative of our findings. The points listed in Table 2 belong to the seesaw models, and those in Table 3 represent the solutions realized in models with $\epsilon=0.05$ and $\epsilon=0.2$, as indicated in the header of Table 3. All of the benchmark points of seesaw scenario presented in Table 2 are compatible with the muon $g-2$ solutions within $2\sigma$, the Planck bound on the LSP neutralino relic density, and the constraint from $\mu\rightarrow e\gamma$. In addition, all these points, including P1, remain outside the contours of the current LHC bounds on the chargino and stau masses.